Find the dimensional formula of ε0.

Grade:11

3 Answers

Aditi Chauhan
9 years ago
Sol. ε base 0 = Coulomb^2/Newton m^2 = l^1 M^-1 L^-3 T^4 ∴ F = kq base 1q base 2/r^2
Apoorva Arora IIT Roorkee
9 years ago
The permittivity of free space is given by
\epsilon _{0}=8.85\times 10^{-12} m^{-3} Kg^{-1}s^{4}A^{2}
so the dimensional formula is
=[M^{-1}L^{-3}T^{4}A^{2}]
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty
Apoorva Arora IIT Roorkee
9 years ago
The permittivity of free space is given by
Description: \epsilon _{0}=8.85\times 10^{-12} m^{-3} Kg^{-1}s^{4}A^{2}
so the dimensional formula is
Description: =[M^{-1}L^{-3}T^{4}A^{2}]
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty